OpenStudy (anonymous):

64x^3+1 can someone factor this and show me the work pleease

OpenStudy (anonymous):

do you know the formula for acube + bcube??

OpenStudy (anonymous):

sum of two cubes (a+b)(a^2-ab+b^2)

OpenStudy (anonymous):

can you walk me through how to do this problem im very lost

OpenStudy (anonymous):

a=4x, b=1

OpenStudy (anonymous):

a+b = 4x+1

OpenStudy (anonymous):

a^2=(4x)^2=16x^2 ab = (4x)(1)=4x b^2=1^2=1

OpenStudy (anonymous):

(a+b)(a^2-ab+b^2) (4x+1)(16x^2-4x+1)

OpenStudy (anonymous):

and is that the answer?

OpenStudy (anonymous):

(4x+1)(16x^2-4x+1) yes, if you multiply that you will get 64x^3 + 1

OpenStudy (anonymous):

thank you can i ask for a few more? please this is all a blur to me.